OpenStudy (anonymous):

Side AB of rectangle ABCD is contained within the line y = 2x - 7. What is the slope of the adjacent side, BC?

OpenStudy (anonymous):

What property do adjacent sides of a rectangle have?

OpenStudy (anonymous):

I dont know what you mean by that?

OpenStudy (anonymous):

At what angle do the adjacent sides meet at?

OpenStudy (anonymous):

a 90 degree angle?

OpenStudy (anonymous):

Good. Now what do you know about the slope of lines which meet at a \(90^{\circ}\) angle?

OpenStudy (anonymous):

that its 2x

OpenStudy (anonymous):

If two lines are perpendicular, then their slopes will multiply to \(-1\). So for example, \(y=3x+1\) and \(y=-\frac{1}{3}-7\) are perpendicular because \((3)\left(-\frac{1}{3}\right)=-1\).

OpenStudy (anonymous):

so the slope is -1/2?

OpenStudy (anonymous):

\(y=-\frac{1}{3}x-7\)*

OpenStudy (anonymous):

Exactly. Excellent.

OpenStudy (anonymous):

Thanks. can you help me with another question. this one is parallel lines

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Which of the following is the equation of a line parallel to 4y = 12x + 5 that passes through (3,7)?

OpenStudy (anonymous):

Do you prefer point-slope form, for example \(y-3=15(x-4)\), or slope-intercept form, for example \(y=13x+2\)?

OpenStudy (anonymous):

The first thing to do, either way, is to figure out the slope of the line they gave you and then think about how the slopes of two parallel lines are related.

OpenStudy (anonymous):

If i divide each side by 4 will i end up like this y=3x+5 but i know that isn't correct.

OpenStudy (anonymous):

Actually, you will have \(y = 3x + \frac{5}{4}\), which has a slop of what?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Now what slope will a parallel line have?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Good, and we know it must pass through \((3, 7)\), so we can form a line from this information. Do you know how?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay, we know that the line has to have the from \(y = mx + b\) and we know \(m=3\). We also know that \((3,7)\) has to be on the line, so it must be a solution to \(y = 3x+b\), whatever \(b\) is (we don't know yet). So we put \((3,7)\) into the equation of our line and solve for \(b\).

OpenStudy (anonymous):

so 7=9 + b?

OpenStudy (anonymous):

Right. So what is \(b\)?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Good. So what is the final equation?

OpenStudy (anonymous):

y=3x-2?

OpenStudy (anonymous):

Yup. :)

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

You're very welcome.